# Bayesian mixture modeling using a mixture of finite mixtures with normalized inverse Gaussian weights

### v1

@@ 2.3 Data augmentation and conditional distribution of $M_{na}$
We illustrate the data augmentation and conditional posterior distribution of $M_{na}$. We introduce the data augmentation in models (2.1) and (2.2). This technique is employed in James et al. (2009) and Argiento and De Iorio (2022). Let $\mathcal{M}_{a}$ and $\mathcal{M}_{na}$ be the allocated and unallocated index sets, respectively. The conditional joint distribution of $\xi=(\xi_{1},\dots,\xi_{M})^{\top}$ and $S=(S_{1},\dots,S_{M})^{\top}$ given $M$ and a label vector $c$ is
$p(S,\tau|M,c)$ $\propto\left(\prod_{i=1}^{n}\frac{S_{c_{i}}}{T}\right)\left(\prod_{m=1}^{M}h(s_{m})p(\tau_{m})\right)$
$=\left(\prod_{m\in\mathcal{M}^{a}}\left(\frac{S_{m}}{T}\right)^{n_{m}}h(s_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}^{na}}h(s_{m})p(\tau_{m})\right),$
where $n_{m}=\#\{i:c_{i}=m\}$. Since this equation consists of categorical distributions, conjugacy with such a distribution is required as $P_{\pi}$ for Gibbs sampling. This restriction is relaxed by using latent a random variable $U_{n}|T\sim\mathrm{Gamma}(n,T)$, where $T=\sum_{m=1}^{M}S_{m}$. In fact,
$p(S,\tau,U_{n}|M,c)\propto u^{n-1}\left(\prod_{m\in\mathcal{M}}e^{-S_{m}u}S_{m}^{n_{m}}h(s_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}^{na}}e^{-S_{m}u}h(s_{m})p(\tau_{m})\right).$
@@
$S_{m}|U_{n},M,c\overset{\mathrm{ind}}{\sim}e^{-S_{m}u}S_{m}^{n_{m}}h(s_{m}),\quad m\in\mathcal{M}_{a},$ (2.4)
$S_{m}|U_{n},M,c\overset{\mathrm{i.i.d}}{\sim}e^{-S_{m}u}h(s_{m}),\quad m\in\mathcal{M}_{na}.$ (2.5)
@@
Under this data augmentation, the conditional distribution of $M_{na}$ is established in Theorem 5.1 in Argiento and De Iorio (2022). This theorem states that if $\mathcal{P}(q_{M},h,p_{0})$ follows a Norm-IFPP, then the posterior distribution of the random measure $P$, given $U_{n}$, is a superposition of a finite point process with fixed points and an IFPP. The IFPP characterizes the process of unallocated jumps, where the discrete probability distribution that serves as its parameter corresponds to the distribution of $M_{na}$. The conditional distribution of $M_{na}$ is given by
$\mathbb{P}(M_{na}=m|\theta,U_{n}=u)\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=1,2,\dots,$ (2.6)
where $\theta=(\theta_{1},\dots,\theta_{n})^{\top}$, $k$ is the number of cluster (unique values of $\theta$) and $\psi$ is a Laplace transform of $h$. We sample $M_{na}$ from (2.6) in the MCMC algorithm. Then, we straightforwardly obtain $M$ by adding $k$ to $M_{na}$.
Remark 2.1.
Although we use the data augmentation $U_{n}|T\sim\text{Gamma}(n,T)$, the distributions (2.4) and (2.5) are generally complex. In addition, the result of Argiento and De Iorio (2022) applies only to $h$ that have a Laplace transform. Thus, the inverse Gaussian distribution is one of the few examples that satisfy both computational and theoretical constraints.
@@ 3 Methodology
In this section, we propose a mixture of finite mixtures with normalized inverse Gaussian weights. Moreover, we develop an efficient posterior sampling algorithm based on Argiento and De Iorio (2022).
@@ 3.1 Mixture of finite mixtures with normalized inverse Gaussian weights
We propose a mixture of finite mixtures with normalized inverse Gaussian weights (denoted by MFM-Inv-Ga), where the notation explicitly specifies $h$ because it is essential for computation. Our proposal model only requires (2.2) to be
$S_{m}|M\overset{\mathrm{ind}}{\sim}\text{Inv-Ga}(\alpha,1),\quad\alpha>0,$
$\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim\text{Norm-Inv-Ga}(\alpha,\dots,\alpha),$
[…]
$h(s)=\frac{\alpha}{\sqrt{2\pi}}s^{-3/2}\exp\left(-\frac{1}{2}\left(\frac{\alpha^{2}}{s}+\beta^{2}s\right)+\beta\alpha\right),$ (3.1)
@@
From Lijoi et al. (2005), the probability density function of the normalized inverse Gaussian distribution with parameters $\alpha_{i}>0,\ i=1,\dots,d$ is given by
@@
where $\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})=\sum_{i=1}^{d-1}\alpha_{i}^{2}/\pi_{i}+\alpha_{d}^{2}\left(1-\sum_{i=1}^{d-1}\pi_{i}\right)^{-1}$ and $K_{m}$ is the modified Bessel functions of the third kind of order $m$ (see, also Ghosal and van der Vaart, 2017). Due to the breakdown of conjugacy with the categorical distribution, constructing an efficient Gibbs sampler is challenging. Because of the complexity of (3.3), the calculation in Miller and Harrison (2018) is intractable: constructing an MCMC algorithm based on the restaurant process by marginalizing out $M$ is extremely challenging. To overcome this difficulty, we use the data augmentation in Subsection 2.3 and the method proposed in Argiento and De Iorio (2022). As a result, it is sufficient to use not (3.3) but (3.1) and (3.8) to construct our MCMC algorithm.
@@ 3.2 Posterior computation
We present a fast and efficient posterior computation algorithm for the proposed method. To this end, we adopt the blocked Gibbs sampling scheme in Argiento and De Iorio (2022). Let $\eta$ be a hyper-parameter of the prior distribution $q_{M}$, and $p(\cdot)$ be a joint prior density function except for $h$ and $q_{M}$. $n_{m}=\#\{i;c_{i}=m\}$ is the size of the $m$th cluster. $\psi(u)$ is the Laplace transform of $h$, which is defined by $\psi(u):=\mathbb{E}[e^{-uS_{m}}]$ for $u\geq 0$.
Algorithm 1 Block Gibbs sampler
@@
  Step 2. Sample $c_{i},for\ i=1,\dots,n,$ the following discrete probability distribution defined by each probabilities
$\mathbb{P}(c_{i}=j|\mathrm{rest})\propto S_{j}f(y_{i}|\tau_{j}),\quad j=1,\dots,M.$
@@
$p(\eta|\mathrm{rest})\propto\Psi(u,k)p(\eta),\quad\Psi(u,k):=\sum_{m=0}^{\infty}\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k).$
@@
$\mathbb{P}(M_{na}=m|\mathrm{rest})\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=0,1,\dots.$
[…]
$p(\tau_{m}|\mathrm{rest})\propto\left\{\prod_{i;\theta_{i}=\tau_{m}}f(y_{i}|\tau_{m})\right\}p(\tau).$
  Step 7. Sample $S_{m_{na}}$, for $\ m_{na}=k+1,\dots,M_{na}+k$from $\mathrm{GIG}(2u+1,\alpha^{2},-1/2).$
[…]
@@ 3.3 Prior distributions for the number of mixture components
Assume that $M-1$ follows a discrete probability distribution with the support $\{0,1,2,\dots\}$. In this paper, we consider $\text{Poisson}(\Lambda)$ ($\Lambda>0$) for $M-1$, because the constraints in Steps 3 and 4 are satisfied. In fact, from Argiento and De Iorio (2022), we have
@@
$p(\Lambda|\mathrm{rest})\propto\psi(u)(k+a_{\Lambda}-1)$ $\mathrm{Ga}(k+a_{\Lambda}+1,1-\psi(u)+b_{\Lambda})$
$+k(b_{\Lambda}+1-\psi(u))\mathrm{Ga}(k+a_{\Lambda},1-\psi(u)+b_{\Lambda}),$
[…]
@@ 3.4.1 Cluster analysis and density estimation
One of the most famous and useful kernels is the (multivariate) normal kernel $N(\cdot|\mu,\sigma^{2})$. In the univariate case, we often use the normal inverse gamma model $N(\mu|m_{0},\sigma^{2}/\eta)\times\text{IG}(\sigma^{2}|c_{0},C_{0})$ as a prior distribution of $\xi=(\mu,\sigma^{2})$, where $\ m_{0}\in\mathbb{R}$, $\eta,\ c_{0},\ C_{0}>0$. The parameter $\eta$ is called a smoothing parameter and plays an important role in density estimation. It is possible to include a hierarchical prior for $\eta$. In the multivariate case, we often employ the normal inverse Wishart model $N_{r}(\mu|m_{0},\Sigma/B_{0})\times\mathrm{W}^{-1}(\Sigma|c_{0},C_{0})$ as a prior distribution of $(\mu,\Sigma)$, where $\ m_{0}\in\mathbb{R}^{r},\ B_{0}>0,\ c_{0}>r-1$ and $C_{0}$ is a positive definite matrix. We will use normal kernels in later numerical experiments in Sections 4.1 and 4.2 for clustering and density estimation. In the context of finite mixture models, various kernels have been proposed. If we have prior information that the data have skewness, the skew normal or skew-t kernel is also useful (see e.g., Frühwirth-Schnatter and Pyne, 2010). The skew-normal and skew-t distributions is easy to handle because these kernels have the scale mixtures of normal representations, except for the degree of freedom of skew-t distribution.

### v2

@@ 2.3 Data augmentation and conditional distribution of $M_{na}$
We illustrate the data augmentation and conditional posterior distribution of $M_{na}$. We introduce the data augmentation in models (2.1) and (2.2). This technique is employed in james2009posterior and argiento2022infinity. Let $\mathcal{M}_{a}$ and $\mathcal{M}_{na}$ be the allocated and unallocated index sets, respectively. The conditional joint distribution of $\tau=(\tau_{1},\dots,\tau_{M})^{\top}$ and $S=(S_{1},\dots,S_{M})^{\top}$ given $M$ and a label vector $c$ is
$p(S,\tau\mid M,c)$ $\propto\left(\prod_{i=1}^{n}\frac{S_{c_{i}}}{T}\right)\left(\prod_{m=1}^{M}h(S_{m})p(\tau_{m})\right)$
$=\left(\prod_{m\in\mathcal{M}_{a}}\left(\frac{S_{m}}{T}\right)^{n_{m}}h(S_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}_{na}}h(S_{m})p(\tau_{m})\right),$
where $n_{m}=\#\{i:c_{i}=m\}$ and $T=\sum_{m=1}^{M}S_{m}$, which depends on $S$. Furthermore, using $U_{n}\mid T\sim\mathrm{Gamma}(n,T)$, we obtain
$p(S,\tau,U_{n}\mid M,c)\propto u^{n-1}\left(\prod_{m\in\mathcal{M}_{a}}e^{-S_{m}U_{n}}S_{m}^{n_{m}}h(S_{m})p(\tau_{m})\right)\left(\prod_{m\in\mathcal{M}_{na}}e^{-S_{m}U_{n}}h(S_{m})p(\tau_{m})\right).$
@@
$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{ind}}{\sim}e^{-S_{m}u}S_{m}^{n_{m}}h(S_{m}),\quad m\in\mathcal{M}_{a},$ (2.5)
$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{i.i.d}}{\sim}e^{-S_{m}u}h(S_{m}),\quad m\in\mathcal{M}_{na}.$ (2.6)
@@
Under this data augmentation, the conditional distribution of $M_{na}$ is established in Theorem 5.1 in argiento2022infinity. This theorem states that if $\mathcal{P}(q_{M},h,p_{0})$ follows a Norm-IFPP, then the posterior distribution of the random measure $P$, given $U_{n}$, is a superposition of a finite point process with fixed points and an IFPP. The IFPP characterizes the process of unallocated jumps, where the discrete probability distribution that serves as its parameter corresponds to the distribution of $M_{na}$. The conditional distribution of $M_{na}$ is given by
$\mathbb{P}(M_{na}=m\mid\theta,U_{n}=u,k)\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=1,2,\dots,$ (2.7)
where $\theta=(\theta_{1},\dots,\theta_{n})^{\top}$, $k$ is the number of clusters (unique values of $\theta$) and $\psi$ is the Laplace transform of $h$. We sample $M_{na}$ from (2.7) in the MCMC algorithm. Then, we straightforwardly obtain $M$ by adding $k$ to $M_{na}$.
@@ 3 Methodology
In this section, we propose a mixture of finite mixtures with normalized inverse Gaussian weights. The posterior inference based on the proposed model can be efficiently implemented using the blocked Gibbs sampling proposed by argiento2022infinity.
@@ 3.1 Mixture of finite mixtures with normalized inverse Gaussian weights
We propose a mixture of finite mixtures with normalized inverse Gaussian weights (denoted by MFM-IGau), where the notation explicitly specifies $h$ because it is essential for computation. Our proposal model only requires (2.2) to be
$S_{m}\mid M\overset{\mathrm{ind}}{\sim}\text{IGau}(\alpha,1),\quad\alpha>0,$
$\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim\text{NIGau}(\alpha,\dots,\alpha),$
where $\text{IGau}(\alpha,\beta)$ is the inverse Gaussian distribution with shape parameter $\alpha>0$ and scale parameter $\beta>0$, and $\text{NIGau}(\alpha_{1},\dots,\alpha_{d})$ is the normalized inverse Gaussian distribution with parameters $\alpha_{i}>0$ for $i=1,\dots,d$. The probability density function of a random variable $S$ according to $\text{IGau}(\alpha,\beta)$ is given by
$h(s)=\frac{\alpha}{\sqrt{2\pi}}s^{-3/2}\exp\left(-\frac{1}{2}\left(\frac{\alpha^{2}}{s}+\beta^{2}s\right)+\beta\alpha\right),\quad s>0,$ (3.1)
@@
Let $\psi_{\text{IGau}}(u)$ be the Laplace transforms of $\text{IGau}(\alpha,1)$. $\psi_{\text{IGau}}(u)$ is the following:
$\psi_{\text{IGau}}(u)=\exp\left(\alpha\left(1-\sqrt{1+2u}\right)\right),\quad u\geq 0.$ (3.3)
Also, cumulant $\kappa_{\mathrm{IGau}}(n,u):=(-1)^{n}\frac{d^{n}}{du^{n}}\psi_{\mathrm{IGau}}(u)$ is
$\kappa_{\mathrm{IGau}}(u;n)=\alpha^{n}\psi_{\text{IGau}}(u)(1+2u)^{-n/2}\frac{K_{n-1/2}\left(\alpha\sqrt{1+2u}\right)}{K_{1/2}\left(\alpha\sqrt{1+2u}\right)},$ (3.4)
where $K_{m}$ is the modified Bessel functions of the second kind of order $m$. The derivation of (3.4) is provided in the Supplementary Material. From lijoi2005hierarchical, the probability density function of the normalized inverse Gaussian distribution with parameters $\alpha_{i}>0,\ i=1,\dots,d$ is given by
@@
where $\mathcal{A}_{d}(\pi_{1},\dots,\pi_{d-1})=\sum_{i=1}^{d-1}\alpha_{i}^{2}/\pi_{i}+\alpha_{d}^{2}\left(1-\sum_{i=1}^{d-1}\pi_{i}\right)^{-1}$.
The normalized inverse Gaussian distribution is less informative and more robust to parameter specification than the Dirichlet distribution. In the context of DMFM, shape parameters such as $\gamma/M$ or $\alpha/M$ become small, and information measures such as entropy and the Gini index of $\mathrm{Dirichlet}(\gamma/M,\dots,\gamma/M)$ also tends to be very small. Therefore, it is desirable to incorporate learning for $\gamma$. In contrast, those of $\mathrm{NIGau}(\alpha/M,\dots,\alpha/M)$ does not decrease as easily, so stable posterior inference can be achieved regardless of the choice of $\alpha$. This fact is confirmed through the numerical experiment in Section 4.1.1. Further analysis of the Gini index is provided in the Supplemental Material. Also, in the context of infinite mixtures, it is known that the prior on the number of clusters induced by the normalized inverse Gaussian process has heavier tails than that induced by the Dirichlet process (lijoi2005hierarchical). It is expected that this robustness carries over to the finite mixture setting as well. These provide the main motivations for using the normalized inverse Gaussian distribution as mixing weights.
Due to the breakdown of conjugacy with the categorical distribution, constructing an efficient Gibbs sampler is challenging. Because of the complexity of (3.5), the calculation in miller2018mixture is intractable: constructing MCMC algorithm based on the restaurant process by marginalizing out $M$ is extremely challenging. To overcome this difficulty, we use the data augmentation technique and the sampling method proposed by argiento2022infinity. As a result, it is no longer necessary to deal directly with the normalized inverse Gaussian distribution (3.5), and the simple MCMC algorithm can be constructed using only the inverse Gaussian distribution (3.1) and its Laplace transform.
@@ 3.2 Posterior computation
In this section, we present a fast and efficient posterior computation algorithm for the proposed method. To this end, we adopt the blocked Gibbs sampling scheme in argiento2022infinity. Moreover, we discuss the blocked Gibbs sampling in detail and explain the computational advantages of using inverse Gaussian weights.
Let $\eta$ and $\rho$ be hyperparameters of $q_{M}$ and $h$, respectively, and $p(\cdot)$ be a joint prior density function except for $h$ and $q_{M}$. $n_{m}=\#\{i;c_{i}=m\}$ is the size of the $m$th cluster. $\psi(u)$ is the Laplace transform of $h$, which is defined by $\psi(u):=\mathbb{E}[e^{-uS_{m}}]$ for $u\geq 0$. $\kappa(u;n)$ is the cumulant of $h$, which is defined by $\kappa(u;n):=(-1)^{n}\frac{d^{n}}{du^{n}}\psi(u)$. Note that $\psi(u)$ and $\kappa(u;n)$ depend on $\rho$.
Algorithm 1 blocked Gibbs sampling (argiento2022infinity)
@@
Step 2. Sample $c_{i}$, for $i=1,\dots,n$ from the following discrete probability distribution defined by each probability
$\mathbb{P}(c_{i}=j\mid\mathrm{rest})\propto S_{j}f(y_{i}\mid\tau_{j}),\quad j=1,\dots,M.$
@@
$p(\eta\mid\mathrm{rest})\propto\Psi(u,k)p(\eta),\quad\Psi(u,k):=\sum_{m=0}^{\infty}\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k).$
Step 3’. Sample the hyperparameter $\rho$ of the $h$ from
$p(\rho\mid\mathrm{rest})\propto\Psi(u,k)\prod_{j=1}^{k}\kappa(u;n_{j})p(\rho).$
@@
$\mathbb{P}(M_{na}=m\mid\mathrm{rest})\propto\frac{(m+k)!}{m!}\psi(u)^{m}q_{M}(m+k),\quad m=0,1,\dots.$
Step 5. Sample $S_{m}$, for $m=1,\dots,k$ from $e^{-S_{m}u}S_{m}^{n_{m}}h(S_{m}).$
Step 6. Sample $\tau_{m}$, for $m=1,\dots,k$ from
$p(\tau_{m}\mid\mathrm{rest})\propto\left\{\prod_{i;\theta_{i}=\tau_{m}}f(y_{i}\mid\tau_{m})\right\}p(\tau).$
Step 7. Sample $S_{m_{na}}$, for $m_{na}=k+1,\dots,M_{na}+k$ from $e^{-S_{m}u}h(S_{m}).$
Step 8. Sample $\tau_{m_{na}}$, for $m_{na}=k+1,\dots,M_{na}+k$ from the prior $p(\tau)$.
We summarize the algorithm in Algorithm 1. The point of this algorithm is the data augmentation through the latent variable such as $U_{n}\mid T\sim\mathrm{Gamma}(n,T)$, where $n$ is the sample size and $T=\sum_{m=1}^{M}S_{m}$. It is important to sample the number of empty components $M_{na}$ from (2.7) in Step 4. This step allows for direct sampling of $M$ by adding $k$ and label variables to be updated as in the finite mixture model with given $M$ in step 2. The update is the same as the telescoping sampling proposed by fruhwirth2021generalized, and the method is more efficient than the classical restaurant process. However, in implementation, it is important to determine $q_{M}$ so that the series $\Psi(u,k)$ in Steps 3 and 4 can be written analytically and random variables can be easily generated from the full conditional distributions of $\eta$ and $M_{na}$. In Step 5 and 7, the allocated and unallocated weights are updated based on (2.5) and (2.6), respectively. The prior $h$ of $S_{m}$ is discussed in detail in the next paragraph. In Steps 7, 8, the assigned unnormalized weights and kernel parameters are updated. When $M_{na}$ sampled in Step 4 is greater than or equal to $1$, Steps 7 and 8 are executed.
To implement the blocked Gibbs sampling, at least two conditions must be satisfied: 1) both the probability density function and the Laplace transform must be available in closed form, 2) each $S_{m}$ can be easily generated from (2.5) and (2.6). Additionally, it is desirable that the cumulants of $h$ can be expressed in closed form when updating $\eta$ or when extending the (static) MFM to the DMFM. When $S_{m}$ follows the $\sigma$-stable distribution, the blocked Gibbs sampling, in general, cannot be used, as the probability density function is available in integral form. Alternatively, a marginal Gibbs sampler based on the restaurant process is required, as employed in argiento2022infinity.
The inverse Gaussian distribution satisfies all of the conditions. The probability distribution, the Laplace transform and the cumulant are given by (3.1), (3.3) and (3.4) respectively. If $h$ is the $\text{IGau}(\alpha,1)$, the full conditional distribution of $S_{m}$ is the generalized inverse Gaussian distribution. From (2.5) and (2.6), we have
$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{ind}}{\sim}\mathrm{GIG}(2u+1,\alpha^{2},n_{m}-1/2),\quad m\in\mathcal{M}_{a},$
$S_{m}\mid U_{n}=u,M=m,c\overset{\mathrm{i.i.d}}{\sim}\mathrm{GIG}(2u+1,\alpha^{2},-1/2),\quad m\in\mathcal{M}_{na},$
where $\mathrm{GIG}(a,b,c)$ is the generalized inverse Gaussian distribution with $a>0,b>0,c\in\mathbb{R}$.
@@ 3.3 Prior distributions for the number of mixture components
Assume that M1M-1 follows a discrete probability distribution with the support $\{0,1,2,\dots\}$. In this paper, we consider $\text{Poisson}(\Lambda)$ ($\Lambda>0$) for M1M-1, because the constraints in Steps 3 and 4 are satisfied. In fact, from argiento2022infinity, we have
@@
$p(\Lambda\mid\mathrm{rest})\propto\psi(u)(k+a_{\Lambda}-1)$ $\mathrm{Gamma}(k+a_{\Lambda}+1,1-\psi(u)+b_{\Lambda})$
$+k(b_{\Lambda}+1-\psi(u))\mathrm{Gamma}(k+a_{\Lambda},1-\psi(u)+b_{\Lambda}),$
$\mathbb{P}(M_{na}=m\mid\mathrm{rest})\propto\Lambda\psi(u)$ $\textrm{Shifted-Poisson}(\Lambda\psi(u),1)+k\mathrm{Poisson}(\Lambda\psi(u)),$
respectively, where $\text{Shifted-Poisson}(\Lambda,t)$ is parallel shifted of the distribution $\text{Poisson}(\Lambda)$ by $t$. Furthermore, if $b_{\Lambda}$ is hierarchically assigned a beta-prime prior $\text{Betaprime}(a_{p},b_{p}),a_{p},b_{p}>0$, then the marginal prior of M1M-1 follows a beta-negative binomial distribution $\text{BNB}(a_{\Lambda},a_{p},b_{p})$. fruhwirth2021generalized recommend using $\text{BNB}(1,4,3)$ under Dirichlet weights. It should be noted, however, that updating $b_{\Lambda}$ requires the Metropolis–Hastings algorithm. Therefore, various discrete distributions can be incorporated into the prior for M1M-1 by using a Poisson-based hierarchical structure.
@@ 3.4.1 Cluster analysis and density estimation
One of the most famous and useful kernels is the (multivariate) normal kernel $N(y\mid\mu,\sigma^{2})$. In the univariate case, we often use the normal inverse gamma model as a prior distribution of $\tau=(\mu,\sigma^{2})$: $\mu\mid\sigma^{2}\sim N(m_{0},\sigma^{2}/\eta),\quad\sigma^{2}\sim\text{IG}(c_{0},C_{0})$, where $m_{0}\in\mathbb{R}$, $\eta,\ c_{0},\ C_{0}>0$. The parameter $\eta$ is called a smoothing parameter and plays an important role in density estimation. It is possible to include a hierarchical prior for $\eta$. In the multivariate case, the hierarchical normal inverse Wishart model: $\mu\sim N_{r}(b_{0},B_{0}),\Sigma^{-1}\mid C\sim W(c_{0},C),C\sim W(g_{0},G_{0})$, where $b_{0}\in\mathbb{R}^{r}$, $B_{0}$ is an $r\times r$ covariance matrix, $c_{0},g_{0}>r-1$, and $G_{0}$ is an $r\times r$ positive definite symmetric matrix. We used normal kernels in later numerical experiments in Sections 4.1 and 4.2 for clustering and density estimation. Of course, other kernels can also be used depending on the purpose (see e.g., fruhwirth2010bayesian).